# FMA-Net: Flow-Guided Dynamic Filtering and Iterative Feature Refinement with Multi-Attention for Joint Video Super-Resolution and Deblurring

### v1

@@ Abstract
We present a joint learning scheme of video super-resolution and deblurring, called VSRDB, to restore clean high-resolution (HR) videos from blurry low-resolution (LR) ones. This joint restoration problem has drawn much less attention compared to single restoration problems. In this paper, we propose a novel flow-guided dynamic filtering (FGDF) and iterative feature refinement with multi-attention (FRMA), which constitutes our VSRDB framework, denoted as FMA-Net. Specifically, our proposed FGDF enables precise estimation of both spatio-temporally-variant degradation and restoration kernels that are aware of motion trajectories through sophisticated motion representation learning. Compared to conventional dynamic filtering, the FGDF enables the FMA-Net to effectively handle large motions into the VSRDB. Additionally, the stacked FRMA blocks trained with our novel temporal anchor (TA) loss, which temporally anchors and sharpens features, refine features in a course-to-fine manner through iterative updates. Extensive experiments demonstrate the superiority of the proposed FMA-Net over state-of-the-art methods in terms of both quantitative and qualitative quality. Codes and pre-trained models are available at: https://kaist-viclab.github.io/fmanet-site.
@@ 1 Introduction
We propose FMA-Net, a novel VSRDB framework based on Flow-Guided Dynamic Filtering (FGDF) and an Iterative Feature Refinement with Multi-Attention (FRMA), to allow for small-to-large motion representation learning with good joint restoration performance (Fig. FMA-Net: Flow-Guided Dynamic Filtering and Iterative Feature Refinement with Multi-Attention for Joint Video Super-Resolution and Deblurring). The key insight of the FGDF is to perform filtering that is aware of motion trajectories rather than sticking to fixed positions, enabling effective handling of large motions with small-sized kernels. Fig. 2 illustrates the concept of our FGDF. The FGDF looks similar to the deformable convolution (DCN) [13] but is different in that it learns position-wise $n\times n$ dynamic filter coefficients, while the DCN learns position-invariant $n\times n$ filter coefficients. Our FMA-Net consists of (i) a degradation learning network that estimates motion-aware spatio-temporally-variant degradation kernels and (ii) a restoration network that utilizes these predicted degradation kernels to restore the blurry LR video. The newly proposed multi-attention, consisting of center-oriented attention and degradation-aware attention, enables the FMA-Net to focus on the target frame and utilize the degradation kernels in a globally adaptive manner for VSRDB. The main contributions of this paper can be summarized as follows:
  • We propose a VSRDB framework, called FMA-Net, that predicts spatio-temporally-variant motion-aware degradation and restoration kernels that are robust to large motions. The FMA-Net utilizes a novel multi-attention to adaptively and globally leverage the predicted degradation kernels for VSRDB.
  • A novel flow-guided dynamic filtering (FGDF) is proposed to effectively handle large motions with relatively small-sized kernels. Extensive ablation studies demonstrate that the degradation kernels obtained by our FGDF are able to handle both small and large motions more precisely than conventional dynamic filtering.
  • We empirically show that the proposed FMA-Net significantly outperforms the recent state-of-the-art (SOTA) methods for video SR and deblurring in objective and subjective qualities on the REDS4, GoPro, and YouTube test datasets under a fair comparison, demonstrating its good generalization ability.
@@ 3.4 Overall Architecture
Table 1: Quantitative comparison on REDS4 for ×$\times 4$ VSRDB. All results are calculated on the RGB channel. Red and blue colors indicate the best and second-best performance, respectively. Runtime is calculated on an LR frame sequence of size $180\times 320$. The superscript $*{}^{*}$ indicates that the model is retrained on the REDS [52] training dataset for VSRDB.
@@ 4.1 Comparisons with State-of-the-Art Methods
To achieve VSRDB, we compare our FMA-Net with the very recent SOTA methods: two single-image SR models (SwinIR [46] and HAT [10]), two single-image deblurring models (Restormer [73] and FFTformer [41]), two VSR models (BasicVSR++ [7] and FTVSR [56]), two video deblurring models (RVRT [48] and GShiftNet [43]), and one VSRDB model (HOFFR [18]). Also, we retrain one single-image model ($*{}^{*}$ [73]) and three video models ($*{}^{*}$ [9], $*{}^{*}$ [43], and $*{}^{*}$ [48]) using our training dataset to perform VSRDB for a fair comparison. It should be noted that $*{}^{*}$ [73] is modified to receive concatenated $T$ frames in the channel dimension for video processing instead of a single frame, and we added a pixel-shuffle [59] block at the end to enable SR.
Table 1 shows the quantitative comparisons for the test set, REDS4. It can be observed in Table 1 that: (i) the sequential approaches of cascading SR and deblurring result in error propagation from previous models, leading to a significant performance drop, and the use of two models also increase memory and runtime costs; (ii) the VSRDB methods consistently demonstrate superior overall performance compared to the sequential cascade approaches, indicating that the two tasks are highly inter-correlated; and (iii) our FMA-Net significantly outperforms all SOTA methods including five joint VSRDB methods in terms of PSNR, SSIM, and tOF. Specifically, our FMA-Net achieves improvements of $1.03$ dB and $1.77$ dB over the SOTA algorithms, $*{}^{*}$ [48] and $*{}^{*}$ [9], respectively. The clip-by-clip analyses for REDS4 and the results of all possible combinations of the sequential cascade approaches can be found in the Supplemental, including demo videos.
@@
Table 2 shows the quantitative comparisons on GoPro [51] and YouTube test sets for joint models trained on REDS [52]. When averaged across both test sets, our FMA-Net achieves a performance boost of $2.08$ dB and $1.93$ dB over $*{}^{*}$ [48] and $*{}^{*}$ [43], respectively. This demonstrates that our FMA-Net has good generalization in addressing spatio-temporal degradation generated from various scenes across diverse devices. Figs. FMA-Net: Flow-Guided Dynamic Filtering and Iterative Feature Refinement with Multi-Attention for Joint Video Super-Resolution and Deblurring(a) and 5 show the visual results on three test sets, showing that the images generated by our FMA-Net are visually sharper than those by other methods.
@@ 5 Conclusion
We propose a novel joint video super-resolution and deblurring framework, called FMA-Net, based on our novel flow-guided dynamic filtering (FGDF) and iterative feature refinement with multi-attention (FRMA). We iteratively update features including self-induced optical flow through stacked FRMA blocks, and predict a flow-mask pair with flow-guided dynamic filters, which enables the network to capture and utilize small-to-large motion information. The FGDF leads to a dramatic performance improvement compared to conventional dynamic filtering. Additionally, the newly proposed temporal anchor (TA) loss facilitates model training by temporally anchoring and sharpening unwarped features. Extensive experiments demonstrate that our FMA-Net achieves SOTA joint performances for diverse datasets with significant margins compared to the most recent SOTA video SR and deblurring methods.
@@ Appendix B Detailed Experimental Results
[…]
@@ B.2 Clip-by-clip Results on REDS4
Table 8 shows the performance of the clip-by-clip results on REDS4 [52] for VSRDB methods and all possible combinations of the sequential cascade approaches. It shows that our FMA-Net exhibits the best performance on all REDS4 clips consisting of realistic and dynamic scenes. In particular, compared to $*{}^{*}$ [48], our FMA-Net achieves PSNR improvement of 0.35 dB in Clip 000, a scene with small motion, improvements of 1.62 dB and 1.58 dB in Clips 011 and 020, scenes with large motion, respectively. This demonstrates the superiority of FMA-Net over existing SOTA methods, especially in scenes with large motion.
@@ C.2 Limitation: Object Rotation
$*{}^{*}$ [73] 26.5 0.081 27.29 / 0.7850 / 2.71
$*{}^{*}$ [43] 13.5 0.185 25.77 / 0.7275 / 2.96
$*{}^{*}$ [9] 7.3 0.072 27.06 / 0.7752 / 2.70
$*{}^{*}$ [48] 12.9 0.680 27.80 / 0.8025 / 2.40
[…]
@@
Table 7: Quantitative comparison on REDS4 for ×$\times 4$ VSRDB. All results are calculated on the RGB channel. Red and blue colors indicate the best and second-best performance, respectively. Runtime is calculated on an LR frame sequence of size $180\times 320$. The superscript $*{}^{*}$ indicates that the model is retrained on the REDS [52] training dataset for VSRDB.
@@
$*{}^{*}$ [73] 26.51 / 0.7551 / 0.47 27.09 / 0.7695 / 3.53 30.03 / 0.8579 / 2.82 25.52 / 0.7573 / 4.04 27.29 / 0.7850 / 2.72
$*{}^{*}$ [43] 24.66 / 0.6730 / 0.93 25.66 / 0.7190 / 3.47 28.05 / 0.7995 / 3.50 24.69 / 0.7187 / 3.93 25.77 / 0.7275 / 2.96
$*{}^{*}$ [9] 25.90 / 0.7234 / 0.57 27.07 / 0.7699 / 3.36 29.67 / 0.8475 / 3.01 25.58 / 0.7601 / 3.86 27.06 / 0.7752 / 2.70
$*{}^{*}$ [48] 26.84 / 0.7764 / 0.38 27.76 / 0.7903 / 2.95 30.66 / 0.8694 / 2.60 25.93 / 0.7740 / 3.65 27.80 / 0.8025 / 2.40
[…]
@@
Table 8: Quantitative comparison on REDS4 for ×$\times 4$ VSRDB. All results are calculated on the RGB channel. Red and blue colors indicate the best and second-best performance, respectively. The superscript $*{}^{*}$ indicates that the model is retrained on the REDS [52] training dataset for VSRDB.

### v2

@@ Abstract
We present a joint learning scheme of video super-resolution and deblurring, called VSRDB, to restore clean high-resolution (HR) videos from blurry low-resolution (LR) ones. This joint restoration problem has drawn much less attention compared to single restoration problems. In this paper, we propose a novel flow-guided dynamic filtering (FGDF) and iterative feature refinement with multi-attention (FRMA), which constitutes our VSRDB framework, denoted as FMA-Net. Specifically, our proposed FGDF enables precise estimation of both spatio-temporally-variant degradation and restoration kernels that are aware of motion trajectories through sophisticated motion representation learning. Compared to conventional dynamic filtering, the FGDF enables the FMA-Net to effectively handle large motions into the VSRDB. Additionally, the stacked FRMA blocks trained with our novel temporal anchor (TA) loss, which temporally anchors and sharpens features, refine features in a coarse-to-fine manner through iterative updates. Extensive experiments demonstrate the superiority of the proposed FMA-Net over state-of-the-art methods in terms of both quantitative and qualitative quality. Codes and pre-trained models are available at: https://kaist-viclab.github.io/fmanet-site.
@@ 1 Introduction
We propose FMA-Net, a novel VSRDB framework based on Flow-Guided Dynamic Filtering (FGDF) and an Iterative Feature Refinement with Multi-Attention (FRMA), to allow for small-to-large motion representation learning with good joint restoration performance. The key insight of the FGDF is to perform filtering that is aware of motion trajectories rather than sticking to fixed positions, enabling effective handling of large motions with small-sized kernels. Fig. 2 illustrates the concept of our FGDF. The FGDF looks similar to the deformable convolution (DCN) [13] but is different in that it learns position-wise $n\times n$ dynamic filter coefficients, while the DCN learns position-invariant $n\times n$ filter coefficients.
Our FMA-Net consists of (i) a degradation learning network that estimates motion-aware spatio-temporally-variant degradation kernels and (ii) a restoration network that utilizes these predicted degradation kernels to restore the blurry LR video. The newly proposed multi-attention, consisting of center-oriented attention and degradation-aware attention, enables the FMA-Net to focus on the target frame and utilize the degradation kernels in a globally adaptive manner for VSRDB. We empirically show that the proposed FMA-Net significantly outperforms the recent state-of-the-art (SOTA) methods for video SR and deblurring in objective and subjective qualities on the REDS4, GoPro, and YouTube test datasets under a fair comparison, demonstrating its good generalization ability.
@@ 3.4 Overall Architecture
Table 1: Quantitative comparison on REDS4 for ×$\times 4$ VSRDB. All results are calculated on the RGB channel. Red and blue colors indicate the best and second-best performance, respectively. Runtime is calculated on an LR frame sequence of size $180\times 320$. The superscript indicates that the model is retrained on the REDS [52] training dataset for VSRDB.
@@ 4.1 Comparisons with State-of-the-Art Methods
To achieve VSRDB, we compare our FMA-Net with the very recent SOTA methods: two single-image SR models (SwinIR [46] and HAT [10]), two single-image deblurring models (Restormer [73] and FFTformer [41]), two VSR models (BasicVSR++ [7] and FTVSR [56]), two video deblurring models (RVRT [48] and GShiftNet [43]), and one VSRDB model (HOFFR [18]). Also, we retrain one single-image model (Restormer [73]) and three video models (BasicVSR++ [9], GShiftNet [43], and RVRT [48]) using our training dataset to perform VSRDB for a fair comparison. It should be noted that Restormer [73] is modified to receive concatenated $T$ frames in the channel dimension for video processing instead of a single frame, and we added a pixel-shuffle [59] block at the end to enable SR.
Table 1 shows the quantitative comparisons for the test set, REDS4. It can be observed in Table 1 that: (i) the sequential approaches of cascading SR and deblurring result in error propagation from previous models, leading to a significant performance drop, and the use of two models also increase memory and runtime costs; (ii) the VSRDB methods consistently demonstrate superior overall performance compared to the sequential cascade approaches, indicating that the two tasks are highly inter-correlated; and (iii) our FMA-Net significantly outperforms all SOTA methods including five joint VSRDB methods in terms of PSNR, SSIM, and tOF. Specifically, our FMA-Net achieves improvements of $1.03$ dB and $1.77$ dB over the SOTA algorithms, RVRT [48] and BasicVSR++ [9], respectively. The clip-by-clip analyses for REDS4 and the results of all possible combinations of the sequential cascade approaches can be found in the Supplemental, including demo videos.
@@
Table 2 shows the quantitative comparisons on GoPro [51] and YouTube test sets for joint models trained on REDS [52]. When averaged across both test sets, our FMA-Net achieves a performance boost of $2.08$ dB and $1.93$ dB over RVRT [48] and GShiftNet [43], respectively. This demonstrates that our FMA-Net has good generalization in addressing spatio-temporal degradation generated from various scenes across diverse devices. Figs. 1(a) and 5 show the visual results on three test sets, showing that the images generated by our FMA-Net are visually sharper than those by other methods.
@@ 5 Conclusion
We propose a novel VSRDB framework, called FMA-Net, based on our novel FGDF and FRMA. We iteratively update features including self-induced optical flow through stacked FRMA blocks, and predict a flow-mask pair with flow-guided dynamic filters, which enables the network to capture and utilize small-to-large motion information. The FGDF leads to a dramatic performance improvement compared to conventional dynamic filtering. Additionally, the newly proposed temporal anchor (TA) loss facilitates model training by temporally anchoring and sharpening unwarped features. Extensive experiments demonstrate that our FMA-Net achieves best performances for diverse datasets with significant margins compared to the recent SOTA methods.
Acknowledgement. This work was supported by the IITP grant funded by the Korea government (MSIT): No. 2021-0-00087, Development of high-quality conversion technology for SD/HD low-quality media and No. RS2022-00144444, Deep Learning Based Visual Representational Learning and Rendering of Static and Dynamic Scenes.
@@ A.6 Visualization of FGDF process
A.7 The Number of FRMA Blocks $M$
Table 7 shows the performance of FMA-Net according to the different numbers of FRMA Blocks $M$. It shows that as $M$ increases, the performance of FMA-Net improves, indicating that the stacked FRMA blocks can effectively update features. Besides Table 7, as can also be seen in Table 1 of the main paper, our smallest FMA-Net variant ($M=1$) even shows superior performance than the previous SOTA methods.
@@ Appendix B Detailed Experimental Results
B.1 FMA-Nets
We first introduce FMA-Nets, a lightweight model of FMA-Net. FMA-Nets is a model that changes the number of FRMA blocks, $M$, from the original $4$ to $2$, with no other modifications. Table 8 compares the quantitative performance of FMA-Nets on REDS4 [52] dataset with one VSRDB method (HOFFR [18]), four retrained SOTA methods (Restormer [73], GShiftNet [43], BasicVSR++ [9] , and RVRT [48] ) for VSRDB on REDS [52], and our FMA-Net. Our FMA-Nets demonstrates the second-best performance, maintaining performance while reducing memory usage and runtime.
@@ B.2 Clip-by-clip Results on REDS4
Table 9 shows the performance of the clip-by-clip results on REDS4 [52] for VSRDB methods and all possible combinations of the sequential cascade approaches. It shows that our FMA-Net exhibits the best performance on all REDS4 clips consisting of realistic and dynamic scenes. In particular, compared to RVRT [48], our FMA-Net achieves PSNR improvement of 0.35 dB in Clip 000, a scene with small motion, improvements of 1.62 dB and 1.58 dB in Clips 011 and 020, scenes with large motion, respectively. This demonstrates the superiority of FMA-Net over existing SOTA methods, especially in scenes with large motion.
@@ C.2 Limitation: Object Rotation
$M$ Params (M) Runtime (s) PSNR $\uparrow$ / SSIM $\uparrow$ / tOF $\downarrow$
1 6.3 0.147 28.07 / 0.8109 / 2.24
2 7.4 0.231 28.46 / 0.8212 / 2.08
4 9.6 0.427 28.83 / 0.8315 / 1.92
Table 7: Ablation study on the number of FRMA blocks $M$.
@@
      Restormer [73]       26.5       0.081       27.29 / 0.7850 / 2.71
      GShiftNet [43]       13.5       0.185       25.77 / 0.7275 / 2.96
      BasicVSR++ [9]       7.3       0.072       27.06 / 0.7752 / 2.70
      RVRT [48]       12.9       0.680       27.80 / 0.8025 / 2.40
      FMA-Nets (Ours)       7.4       0.231       28.46 / 0.8212 / 2.08
@@
Table 8: Quantitative comparison on REDS4 for ×$\times 4$ VSRDB. All results are calculated on the RGB channel. Red and blue colors indicate the best and second-best performance, respectively. Runtime is calculated on an LR frame sequence of size $180\times 320$. The superscript indicates that the model is retrained on the REDS [52] training dataset for VSRDB.
@@
Restormer [73] 26.51 / 0.7551 / 0.47 27.09 / 0.7695 / 3.53 30.03 / 0.8579 / 2.82 25.52 / 0.7573 / 4.04 27.29 / 0.7850 / 2.72
GShiftNet [43] 24.66 / 0.6730 / 0.93 25.66 / 0.7190 / 3.47 28.05 / 0.7995 / 3.50 24.69 / 0.7187 / 3.93 25.77 / 0.7275 / 2.96
BasicVSR++ [9] 25.90 / 0.7234 / 0.57 27.07 / 0.7699 / 3.36 29.67 / 0.8475 / 3.01 25.58 / 0.7601 / 3.86 27.06 / 0.7752 / 2.70
RVRT [48] 26.84 / 0.7764 / 0.38 27.76 / 0.7903 / 2.95 30.66 / 0.8694 / 2.60 25.93 / 0.7740 / 3.65 27.80 / 0.8025 / 2.40
FMA-Nets (Ours) 27.08 / 0.7852 / 0.33 28.73 / 0.8164 / 2.46 30.98 / 0.8745 / 2.42 27.03 / 0.8089 / 3.10 28.46 / 0.8212 / 2.08
@@
Table 9: Quantitative comparison on REDS4 for ×$\times 4$ VSRDB. All results are calculated on the RGB channel. Red and blue colors indicate the best and second-best performance, respectively. The superscript indicates that the model is retrained on the REDS [52] training dataset for VSRDB.